5. Suppose a firm offers two software products: Wordy -- a document creator, and Excello, a spreadsheet app. They know there are two types of consumers overall, English types and Quant types, each with a different valuation (demand) for each product. Use the following table to answer questions below.

Assume the cost of producing Wordy is a constant MC = $100, the cost of producing Excello is a constant MC = $150 and that each consumer will purchase each good as long as the price is less than or equal to value.

Consumer values are entries in the following table, by type:

Writing/English types

Quantitative Types

Wordy

$500

$200

Excello

$400

$650

  • (a) Suppose the monopolist only sold the goods separately. What price will the monopolist charge for Wordy to maximize profits. Explain. (hint: assume that there are only 2 consumers).
  • (b) Suppose the monopolist only sold the goods separately. What price will the monopolist charge for Excello to maximize profits?
  • (c) What is the total profit to the monopolist from selling the goods separately?
  • (d) What is a better pricing strategy for the monopolist? What is the resulting profit?
Answer #1

QA) for Wordy

If P = 200, both type will buy the good

So π = P*Q - TC

= 200*2 - 100*2 = 400-200 = 200

.

If P =500, then only English Types will buy

So , π = 500-100 = 400

higher π is when P = 500 is charged

so for Wordy, Charge = $ 500

.

b) for Excello

If P = 400, both type will buy

So π = 400*2 -150*2 = 800-300 = 500

.

If P = 650, only Quant type will buy

π = 650-150. = 500

As same profit when P = 400 or P = 650

so charge either P = 400 or P = $ 650

c) total profit = 400+500

= $ 900

d) if two goods are sold jointly

cost of bundle = 150+100 = 250

Then if P = 900 ( sum of willingness of English type, for both goods)

Then only English type will buy

π = 900-(150+100)

= 650

If P = 850

Both type will buy

π = 2*850 - 2*250 = 1200

thus higher profit , when two goods are sold together

so sell two goods as bundle , at price of bundle = 850

with π = 1200